# Zero-Inflated Tweedie Boosted Trees with CatBoost for Insurance Loss Analytics

### v1

@@ 4.2 Scenario 2: Functionally related
We introduce Scenario 2 as a novel methodology wherein the parameter $q$ is functionally linked to $\mu$. This innovative approach, absent from existing literature, acknowledges the relationship between the inflation parameter $q$ and the mean parameter $\mu$. Given our limited prior understanding of this relationship, we propose a refinemen of the parameterization, initially introduced by Lambert, (1992) for zero-inflated Poisson Generalized Linear Models (GLM).
Our proposed parameterization is depicted by the following equations:
@@
From (23), we derive:
@@
Equation (24) reveals that for $\gamma>0$, the inflation probability diminishes as $\mu$ increases. Guided by this assumption, we define the loss function as follows:
@@
where $\mu=Ee^{W_{T}}$. The gradient and Hessian of the loss function with respect to $W_{t}$ are:
$g_{t+1}=\partial_{W_{t}}\ell(y,W_{t})=\begin{cases}-\displaystyle\frac{\zeta\cdot\eta}{1+\zeta}+\kappa,&\quad y=0\\[12.91663pt] -\gamma+\kappa-\displaystyle\frac{1}{\phi}y\mu^{1-p}+\displaystyle\frac{1}{\phi}\mu^{2-p},&\quad y>0,\\ \end{cases}$ (26)
$h_{t+1}=\partial^{2}_{W_{t}}\ell(y,W_{t})=\begin{cases}\displaystyle\frac{\left(\displaystyle\frac{2-p}{\phi}\mu^{2-p}-\eta^{2}\right)\cdot\zeta\cdot(1+\zeta)+\zeta^{2}\cdot\eta^{2}}{(1+\zeta)^{2}}+\kappa^{2}\cdot\mu^{-\gamma},&\quad y=0\\[12.91663pt] \kappa^{2}\cdot\mu^{-\gamma}-\displaystyle\frac{(1-p)y}{\phi}\mu^{1-p}+\displaystyle\frac{(2-p)}{\phi}\mu^{2-p},&\quad y>0,\\ \end{cases}$ (27)
@@ 4.3 Adjustment of Compositional Data
Compositional data is characterized by multiple non-negative features that sum up to a constant, typically 100% or 1. Due to the inherent statistical dependence among these features, transformations are often necessary to map the data onto the real Euclidean space. This transformation facilitates the application of traditional statistical methodologies (Aitchison, (1994)). Among the notable transformations are the logratio methods, which include the centered logratio transformation (CLR), the additive logratio transformation (ALR), and the isometric logratio transformation (ILR).
@@
The matrix $R$ is referred to as the contrast matrix (Pawlowsky-Glahn et al., (2015)). Although these logratio transformation methods are equivalent up to linear transformations, the ALR does not preserve distances between data points in the original space. While CLR retains such distances, it can lead to singular covariance matrices. On the contrary, the ILR overcomes these issues but results in transformed features that are complex combinations of the original features, making interpretation additionally challenging.
@@ 5.1.1 Deviance and Mean Absolute Deviation
The unit deviance for the $i$th observation in zero-inflated Tweedie models can be determined by the equation below:
@@ 5.2 Synthetic Telematics Data
The synthetic telematics dataset developed by So et al., (2021) used in this study is publicly accessible at http://www2.math.uconn.edu/~valdez/data.html. This dataset was meticulously generated using Synthetic Minority Over-sampling Technique (SMOTE), coupled with a feedforward Neural Network (NN), and trained based on authentic data from a Canadian insurance provider. It consists of 100,000 data samples, with only 2,698 policies reporting at least one claim. This notable occurrence of zero-inflation within the dataset highlights a distinct characteristic of the data. The dataset is structured around 52 distinct features categorized into Traditional data (e.g., exposure, insured age, and gender), Telematics data (e.g., total miles driven, harsh braking, and harsh acceleration), and Claims Data (encompassing claim counts and aggregate claim amounts). Evidently, the aggregate claim amount models were rigorously developed utilizing exposure and 49 features, with the response variable being the claim amounts. Among these 49 features, five are categorical variables, each having varying numbers of categories with 2, 2, 4, 2, and 55 categories, respectively. A noteworthy attribute of this dataset is the presence of a few categorical attributes. This suggests the need for the selection of models capable of effectively handling such features to ensure optimal results and accurate data fitting. To this end, the CatBoost library was employed to implement boosted tree algorithms, which offers the necessary capabilities to handle these intricate features. For a detailed overview of the dataset and its attributes, readers are directed to Table 7 in Appendix B, where a comprehensive description of the variables in the dataset can be found. Figure 2 provides an overview of the distribution of the aggregate claim amounts for the synthetic telematics data. The graph shows a point mass at zero and a histogram of the positive aggregate claim amounts.
Figure 2: Distribution of Aggregate Claim Amounts for the Synthetic Telematics Data
@@ 6 Concluding Remarks
Broadly speaking, for short-term insurance contracts, it is not uncommon to encounter claim portfolios that exhibit a large mass at zero and a right-skewed distribution for positive claims. This distribution is characterized by a prominent peak at zero, indicating a high likelihood of no claims being made. The adoption of “zero-inflated” models, such as the zero-inflated Poisson, negative binomial, and Tweedie regression models, has become more prevalent in practice. While GLM has traditionally been a favored choice in actuarial studies, its linear structure can pose limitations, especially when confronted with nonlinearities and intricate interactions inherent in datasets such as auto telematics data, which captures diverse driving behaviors. Consequently, there is a growing preference for machine learning approaches to better model these complex interactions. Among these approaches, gradient boosting techniques have emerged as powerful tools for constructing predictive models for insurance claims. See Zhou et al., (2022) and and Meng et al., (2022).
In this paper, we integrated the advantages of applying a zero-inflated Tweedie loss function within a gradient boosting algorithm. To achieve this, we employed several adjustments. Specifically, we reparameterized the zero-inflated Tweedie loss function to enable us to express the inflation probability $q$ as a function of $\mu$. This reparameterization resulted in a singular model, which enabled us to optimally utilize the CatBoost libraries. This approach enhances interpretation and facilitates better model comparison by implementing various performance metrics. CatBoost excels at iteratively transforming weak learners into strong learners. Among existing boosting libraries, So, (2024) recommended the use of CatBoost for its efficient implementation in insurance loss estimation.
@@ 1 Initialize trees: $w^{mean}_{0}=0$, $w^{prob}_{0}=0$, $\phi=1$ ;
4       Fit the tree $w^{mean}_{t}$ minimizing regularized loss function(10);
5       Update $\ W^{mean}_{t}=W^{mean}_{t-1}+\alpha w^{mean}_{t}$;

### v2

@@ 4.2 Scenario 2: Functionally related
We introduce Scenario 2 as a novel methodology wherein the parameter $q$ is functionally linked to $\mu$. This innovative approach, absent from existing literature, acknowledges the relationship between the inflation parameter $q$ and the mean parameter $\mu$. Given our limited prior understanding of this relationship, we propose a refinement, within the boosting algorithm, of the parameterization initially introduced by Lambert, (1992) for zero-inflated Poisson Generalized Linear Models (GLM).
Inspired by this work, our proposed parameterization allows us to relate the canonical link functions of the zero component, which is the $\text{logit}(q)$, and the continuous component, which is the $\ln\mu$, as depicted in the following equations:
@@
From (23), we are able to derive the relationship between $q$ and $\mu$:
@@
The parameter $\gamma$ is estimated from the data. Equation (24) reveals that for $\gamma>0$, the inflation probability diminishes as $\mu$ increases. Guided by this assumption, we define the loss function as follows:
@@
where $\mu=Ee^{W_{T}(\bm{x})}$. The gradient and Hessian of the loss function with respect to $W_{t}(\bm{x})$ are:
$g_{t+1}=\partial_{W_{t}(\bm{x})}\ell(y,W_{t}(\bm{x}))=\begin{cases}-\displaystyle\frac{\zeta\cdot\eta}{1+\zeta}+\kappa,&\quad y=0\\[12.91663pt] -\gamma+\kappa-\displaystyle\frac{1}{\phi}y\mu^{1-p}+\displaystyle\frac{1}{\phi}\mu^{2-p},&\quad y>0,\\ \end{cases}$ (26)
$h_{t+1}=\partial^{2}_{W_{t}(\bm{x})}\ell(y,W_{t}(\bm{x}))=\begin{cases}\displaystyle\frac{\left(\displaystyle\frac{2-p}{\phi}\mu^{2-p}-\eta^{2}\right)\cdot\zeta\cdot(1+\zeta)+\zeta^{2}\cdot\eta^{2}}{(1+\zeta)^{2}}+\kappa^{2}\cdot\mu^{-\gamma},&\quad y=0\\[12.91663pt] \kappa^{2}\cdot\mu^{-\gamma}-\displaystyle\frac{(1-p)y}{\phi}\mu^{1-p}+\displaystyle\frac{(2-p)}{\phi}\mu^{2-p},&\quad y>0,\\ \end{cases}$ (27)
@@ 4.3 Adjustment of Compositional Data
Compositional data is characterized by multiple non-negative features that sum up to a constant, typically 100% or 1. Due to the inherent statistical dependence among these features, transformations are often necessary to map the data onto the real Euclidean space. This transformation facilitates the application of traditional statistical methodologies (Aitchison, (1982)). Among the notable transformations are the logratio methods, which include the centered logratio transformation (CLR), the additive logratio transformation (ALR), and the isometric logratio transformation (ILR). See also Aitchison, (1994).
@@
The matrix $R$ is referred to as the contrast matrix (Pawlowsky-Glahn et al., (2015)). Although these logratio transformation methods are equivalent up to linear transformations, the ALR does not preserve distances between data points in the original space. While CLR retains such distances, it can lead to a determinant of zero. On the contrary, the ILR overcomes these issues but results in transformed features that are complex combinations of the original features, making interpretation additionally challenging.
@@ 5.1.1 Deviance and Mean Absolute Deviation
The unit deviance for the $i$-th observation in zero-inflated Tweedie models can be determined by the equation below:
@@ 5.2 Synthetic Telematics Data
The synthetic telematics dataset developed by So et al., (2021) used in this study is publicly accessible at http://www2.math.uconn.edu/~valdez/data.html. This dataset was meticulously generated using Synthetic Minority Over-sampling Technique (SMOTE), coupled with a feedforward Neural Network (NN), and trained based on authentic data from a Canadian insurance provider. It consists of 100,000 data samples, with only 2,698 policies reporting at least one claim. This notable occurrence of zero-inflation within the dataset highlights a distinct characteristic of the data. The dataset is structured around 52 distinct features categorized into Traditional data (e.g., exposure, insured age, and gender), Telematics data (e.g., total miles driven, harsh braking, and harsh acceleration), and Claims Data (encompassing claim counts and aggregate claim amounts). Evidently, the aggregate claim amount models were rigorously developed utilizing exposure and 49 features, with the response variable being the claim amounts. Among these 49 features, five are categorical variables, each having varying numbers of categories with 2, 2, 4, 2, and 55 categories, respectively. A noteworthy attribute of this dataset is the presence of a few categorical attributes. This suggests the need for the selection of models capable of effectively handling such features to ensure optimal results and accurate data fitting. To this end, the CatBoost library was employed to implement boosted tree algorithms, which offers the necessary capabilities to handle these intricate features. For a detailed overview of the dataset and its attributes, readers are directed to Table 7 in Appendix B, where a comprehensive description of the variables in the dataset can be found. Figure 2 provides an overview of the distribution of the aggregate claim amounts for the synthetic telematics data. The graph shows a point mass at zero and a histogram of the positive aggregate claim amounts. For readers who want to know more about this data, please see So et al., (2021).
Figure 2: Distribution of Aggregate Claim Amounts, on a log scale, for the Synthetic Telematics Data
@@ 6 Concluding Remarks
Broadly speaking, for short-term insurance contracts, it is not uncommon to encounter claim portfolios that exhibit a large mass at zero and a right-skewed distribution for positive claims. This distribution is characterized by a prominent peak at zero, indicating a high likelihood of no claims being made. The adoption of “zero-inflated” models, such as the zero-inflated Poisson, negative binomial, and Tweedie regression models, has become more prevalent in practice. While GLM has traditionally been a favored choice in actuarial studies, its linear structure can pose limitations, especially when confronted with nonlinearities and intricate interactions inherent in datasets such as auto telematics data, which captures diverse driving behaviors. Consequently, there is a growing preference for machine learning approaches to better model these complex interactions. Among these approaches, gradient boosting techniques have emerged as powerful tools for constructing predictive models for insurance claims. See Zhou et al., (2022) and Meng et al., (2022).
In this paper, we integrated the advantages of applying a zero-inflated Tweedie loss function within a gradient boosting algorithm. To achieve this, we employed several adjustments. Specifically, we reparameterized the zero-inflated Tweedie loss function to enable us to express the inflation probability $q$ as a function of $\mu$. This reparameterization resulted in a singular model, which enabled us to optimally utilize the CatBoost libraries. This approach enhances interpretation and computational efficiency. CatBoost excels at iteratively transforming weak learners into strong learners. Among existing boosting libraries, So, (2024) recommended the use of CatBoost for its efficient implementation in insurance loss estimation.
@@ 1 Initialize trees: $w^{mean}_{0}=0$, $w^{prob}_{0}=0$, $\phi=1$ ;
4       Fit the tree $W^{mean}_{T}(\bm{x})$ minimizing regularized loss function(10);
5       Update $\ W^{mean}_{T}(\bm{x})=W^{mean}_{t-1}(\bm{x})+\alpha W^{mean}_{T}(\bm{x})$;